# Word Alignment as Preference for Machine Translation

### v1

@@ Abstract
The problem of hallucination and omission, a long-standing problem in machine translation (MT), is more pronounced when a large language model (LLM) is used in MT because an LLM itself is susceptible to these phenomena. In this work, we mitigate the problem in an LLM-based MT model by guiding it to better word alignment. We first study the correlation between word alignment and the phenomena of hallucination and omission in MT. Then we propose to utilize word alignment as preference to optimize the LLM-based MT model. The preference data are constructed by selecting chosen and rejected translations from multiple MT tools. Subsequently, direct preference optimization is used to optimize the LLM-based model towards the preference signal. Given the absence of evaluators specifically designed for hallucination and omission in MT, we further propose selecting hard instances and utilizing GPT-4 to directly evaluate the performance of the models in mitigating these issues. We verify the rationality of these designed evaluation methods by experiments, followed by extensive results demonstrating the effectiveness of word alignment-based preference optimization to mitigate hallucination and omission.
@@ 1 Introduction
Large language models (LLMs) have been evolving rapidly and showing predominant performance in many natural language processing (NLP) tasks (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). However, in machine translation (MT), the use of a decoder-only LLM is still limited due to issues such as model size (Xu et al., 2024a) and low-resource languages (Hendy et al., 2023). Conventional encoder-decoder MT models trained on parallel corpora still dominate in practice (Costa-jussà et al., 2022). One of the primary concerns of applying an LLM to MT is reliability. Although it does not happen frequently, an LLM is known to hallucinate (Dhuliawala et al., 2023; Zhang et al., 2023a; Bang et al., 2023) as it is pre-trained to predict the next token in very large-scale raw texts. Specifically in MT, LLM-based translation systems therefore could have the phenomena of hallucination and omission, which is also a long-term challenge in the field of MT (Vamvas and Sennrich, 2022), known as over- and under-translation. In this work, we attempt to mitigate the hallucination and omission in LLM-based MT to improve its practicality.
(a) Coverage score distribution of different omission degree.
(b) Coverage score distribution of different omission degree.
Figure 1: The coverage score distribution of degrees of hallucination and omission. The coverage scores are predicted by word alignment model Wu et al. (2023). The annotation of hallucination and omission is from HalOmi benchmark Dale et al. (2023b). Details about the dataset and word alignment model can be found in §4.2.
@@
Consequently, we propose Word Alignment Preference (WAP) that utilizes word alignment as a signal to optimize LLM-based MT models. WAP consists of three steps: diverse translation collection, preference data construction, and preference optimization. Specifically, we collect diverse translations with multiple existing translation tools, select chosen and rejected examples with the word aligner (Wu et al., 2023), and optimize the model on preference data using direct preference optimization (DPO) (Rafailov et al., 2024).
Furthermore, the evaluation of hallucination and omission is challenging, and there is no existing evaluator specifically designed for this. Improving the BLEU and COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency. In addition, hallucination is relatively infrequent, although very severe once happens. Hence, to effectively evaluate it, we design extensive experiments that include testing on instances that potentially have the problem of hallucination and omission, and using GPT-4 as the evaluator with comprehensive analysis. Experimental analysis demonstrates the effectiveness of WAP in mitigating hallucination and omission in MT.
@@
  • As there is no benchmark particularly for evaluating the performance of MT models on hallucination and omission. We design various experiments, including selecting hard instances and utilizing GPT-4 as the evaluator in §4.3. The effectiveness of the evaluation, as well as the proposed WAP has been validated through extensive experiments and analysis in §5
@@ 2 Related work
2.1 Hallucination and omission in machine translation
@@
The problem of hallucination and omission can also be termed an over- and under-translation. Hallucinations in machine translation are cases in which the model generates output that is partially or completely unrelated to the source sentence, while omissions are translations that do not include some of the input information (Dale et al., 2023b). Dale et al. (2023a) explore methods that leverage the internal workings of models and external tools, such as cross-lingual sentence similarity and natural language inference models, to detect and mitigate hallucinations in machine translation. HalOmi (Dale et al., 2023b) introduces an annotated dataset specifically designed to detect hallucinations and omissions, covering 18 translation directions. In Figure 1 and §4.3 we use HalOmi as a reference to assess how these two phenomena correlate to the coverage output of the GPT-4 evaluator and the word aligner, respectively.
2.2 Preference tuning for LLMs
@@
LLMs are capable of completing tasks in the zero-shot or few-shot manner (Radford et al., 2019; Brown et al., 2020). In addition, performance in downstream tasks can also be enhanced by fine-tuning them with instruction datasets (Wei et al., 2022; Chung et al., 2024). However, acquiring instruction datasets is costly, while obtaining preferences for LLM responses is relatively easier (Rafailov et al., 2024). Consequently, numerous studies have fine-tuned the instruction-tuned models on preference data to better align with the feedback from human or advanced models, thus improving their performance on corresponding tasks (Ouyang et al., 2022; Rafailov et al., 2024; Xu et al., 2024b). InstructGPT (Ouyang et al., 2022) aligns language models with human intentions through a two-stage process: supervised instruction fine-tuning and reinforcement learning from human feedback using proximal policy optimization (Schulman et al., 2017). DPO (Rafailov et al., 2024) directly optimize LLM with preference data. It further simplifies the preference-tuning process by removing an extra reward model. We utilize DPO in this work due to the ease of use and effectiveness. A contemporaneous preference-based MT model ALMA-R (Xu et al., 2024b), built on the foundations of DPO, introduces contrastive preference optimization to fine-tune LLMs specifically using reference-free MT metrics and human annotation as preference. ALMA-R focuses on improving general LLM-based MT while we attempt to mitigate the hallucination and omission in MT, and our preference data are entirely made automatically, which also draws the difference between ALMA-R and our work.
2.3 Word alignment
Word-level information is useful in many NLP tasks such as language pretraining Chi et al. (2021); Wu et al. (2021) and cross-lingual sentence embedding Zhang et al. (2023b); Li et al. (2023); Miao et al. (2024). In particular, word alignment plays an important role in MT Bahdanau et al. (2015); Tu et al. (2016). Recently, word aligners based on pre-trained language models, such as SimAlign Jalili Sabet et al. (2020), AWESoME Dou and Neubig (2021) and SpanAlign Nagata et al. (2020); Chousa et al. (2020), have significantly outperformed previous word aligners based on statistical machine translation, such as Giza++ Och and Ney (2003) and FastAlign Dyer et al. (2013). SimAlign is an unsupervised approach based on the similarity of contextualized word embeddings. AWESoME and SpanAlign are supervised models that are trained on parallel corpora and manual word alignment datasets. WSPAlign (Wu et al., 2023) is a weakly-supervised approach trained on large-scale automatically collected data. WSPAlign is the state-of-the-art word aligner, and hence we use it in this work. In addition, Bahdanau et al. (2015) introduces the “align and transaltion” using attention, which also inspires us to utilize the external word aligner as preference for MT.
@@ 2.3 Word alignment
Figure 2: The illustration of WAP framework. The source text is first translated by multiple MT tools, including human translation. An external word aligner is then utilized to predict the coverage score for each translation. Next, translation with the highest and lowest coverage score are selected as preference pairs for the final preference optimization.
@@ 3.1 Gathering translation candidates
We start with the parallel training data in ALMA (Xu et al., 2024a). This parallel data encompasses five language pairs with human translations in both directions: $cs\leftrightarrow en$, $de\leftrightarrow en$, $is\leftrightarrow en$, $zh\leftrightarrow en$ and $ru\leftrightarrow en$. We employ ISO 639 language codes [^1] https://en.wikipedia.org/wiki/List_of_ISO_639_language_codes to denote languages. Specifically, “$cs$” corresponds to Czech, “$de$” to German, “$is$” to Icelandic, “$zh$” to Chinese and “$ru$” and “$en$” to Russian and English, respectively. To generate the translations we require, this dataset is translated in both directions using two well-known MT tools, including DeepL [^2] https://www.deepl.com/en/translator and ChatGPT (gpt-3.5-turbo-0613) [^3] https://openai.com/product. The prompt for ChatGPT that we utilize to translate sentences is shown in Figure 3. The original human-written translation in the training set is also utilized. In particular, Icelandic ($is$) is not supported by the DeepL API, therefore, we use the Google Translate API [^4] https://cloud.google.com/translate/docs/basic/translate-text-basic as an alternative.
Figure 3: The prompt of ChatGPT that we use to translate sentences.
@@ 3.2 Selecting chosen and rejected translation
After obtaining the translation candidates $(y^{1},...,y^{K})$, we use a state-of-the-art public word aligner, namely WSPAlign [^5] https://github.com/qiyuw/WSPAlign, to automatically annotate the degree of coverage for each translation. We follow the usage setting in the original paper of WSPAlign (Wu et al., 2023). In particular, WSPAlign performs a bidirectional alignment and uses a threshold to filter out low-confident alignment of word pairs. Then, the ratio of the source words, that are aligned with at least one word, in the translation is taken as the coverage score, which will be used for the following preference annotation. The whole process predicting the coverage score is notated as $\mathrm{C}(\cdot,\cdot)$. Formally, the coverage score for a translation $y^{k}$ can be calculated by $\mathrm{C}(x,y^{k})\in[0.0,100.0]$. Subsequently, the preferred translation and the rejected translation are selected by the following criteria:
@@ 3.3 Filtering
Note that the whole pipeline of constructing the preference data is automatic, and existing MT and word alignment models are not perfect. Even for human-annotated translation, the quality of it is also an issue that cannot be ignored (Xu et al., 2024b), and may affect the performance of the model trained on it. Hence, noises are inevitable in both the translated texts and the preference choices. On the other hand, the MT tools we choose generally have good performance, it could happen that the generated translations are not diverse enough, leading to the preference signal being disrupted. To improve the quality of the constructed preference datasets as much as possible, multiple strategies are applied to filter out potential bad training instances:
@@
  • Remove the instance where the chosen and rejected translations are too semantically similar. Sentence embedding is a widely used technique to calculate pairwise sentence similarity with low computation cost Gao et al. (2021); Wu et al. (2022); Zhao et al. (2024). LaBSE Feng et al. (2022) [^6] https://huggingface.co/sentence-transformers/LaBSE is used in our experiments. We notate it as $\mathrm{LB}(\cdot)$. The similarity threshold is empirically set as 0.9, i.e. $(x,y^{w},y^{l})$ is excluded from the dataset if $\mathrm{sim}(\mathrm{LB}(y^{w}),\mathrm{LB}(y^{w}))>0.9$, where $\mathrm{sim}(\cdot,\cdot)\in[0.0,1.0]$ is cosine similarity.
@@
3.4 Details of dataset
Figure 4 presents the varying proportions of “chosen” and “rejected” preference pairs from three sources: ChatGPT, DeepL, and Human. The figure indicates that the majority of the “chosen” translations originate from ChatGPT, while a significant portion of human-written translations are “rejected”. This observation supports the conclusion that human-written translations can also exhibit quality issues, as discussed in ALMA-R (Xu et al., 2024b). Examples in our constructed preference dataset are presented in §A.1.
@@ 3.5 Optimization LLM-based MT model
The final step is to optimize the LLM-based MT model on our preference data. Direct preference optimization (DPO) (Rafailov et al., 2024) is a simple but effective approach that directly optimizes the preference model on a pre-constructed static dataset. DPO has been applied to optimize LLM in preference data (Tunstall et al., 2023; Xu et al., 2024b) recently. We also utilize DPO as an optimization approach. Formally, the training objective is as follows,
@@
where $\sigma$ is the sigmoid function, $\pi$ is the model to optimize and $\pi_{ref}$ is the reference model. We use ALMA-13B [^8] https://github.com/fe1ixxu/ALMA as our base model, i.e., the starting point of $\pi$, in the experiments. ALMA-13B is also used as a reference model $\pi_{ref}$, but note that $\pi_{ref}$ will not be updated during training.
Figure 5: Prompt to calculate the coverage score.
@@
Figure 6: Comparison of WAP and baseline in hard and easy instances. $N$ instances with the lowest COMET score by the baseline are selected from the test set as hard instances, and the remaining are easy instances. Results when $N=100$, $200$ and $500$ are presented. Refer to §B for the full numeric results of the entire test.
@@ 4.2 Baselines and evaluation datasets
We choose ALMA-13B [^11] https://huggingface.co/haoranxu/ALMA-13B as the baseline for all experiments in this paper, as well as the starting point of optimization. ALMA (Xu et al., 2024a) was trained from Llama Touvron et al. (2023) in two steps: initial fine-tuning on monolingual data and subsequent fine-tuning on a small set of high-quality parallel data.
For fairly studying the effect of word alignment preference, we use the data used in the supervised fine-tuning in ALMA as the source dataset to construct our preference data in §3. Specifically, the source data was collected from WMT’17 (Bojar et al., 2017) to WMT’20 (Barrault et al., 2020), in addition to the development and text dataset from Flores-200 (Costa-jussà et al., 2022). After filtering, we finally make 20,074 and 2,226 preference triplets for training and development, respectively. For evaluation, the test set is from WMT22, except that $is\leftrightarrow en$ is from WMT21. The remaining data from WMT21 (except $is\leftrightarrow en$) is used as the development set. Specifically, 3485, 4021, 2000, 3912, 4053 examples are included in the test set for $cs\leftrightarrow en$, $de\leftrightarrow en$, $is\leftrightarrow en$, $zh\leftrightarrow en$, and $ru\leftrightarrow en$, respectively.
@@
In particular, we want to validate whether our proposed method is capable of mitigating hallucination and omission in MT. Hence, we also utilize HalOmi (Dale et al., 2023b) in the experiments. HalOmi is an evaluation benchmark for the detection of hallucination and omission in MT. It contains fine-grained sentence-level and token-level annotations of full and partial hallucinations and omissions that cover 18 language directions. Each instance in the data set was annotated in “No hallucination and omission”, “Small hallucination and omission”, “Partial hallucination and omission” or “Full hallucination and omission” by humans. In this paper, we use it to test the performance of GPT-4 as an evaluator. Details are in §4.3.

### v2

@@ Abstract
The problem of hallucination and omission, a long-standing problem in machine translation (MT), is more pronounced when a large language model (LLM) is used in MT because an LLM itself is susceptible to these phenomena. In this work, we mitigate the problem in an LLM-based MT model by guiding it to better word alignment. We first study the correlation between word alignment and the phenomena of hallucination and omission in MT. Then we propose to utilize word alignment as preference to optimize the LLM-based MT model. The preference data are constructed by selecting chosen and rejected translations from multiple MT tools. Subsequently, direct preference optimization is used to optimize the LLM-based model towards the preference signal. Given the absence of evaluators specifically designed for hallucination and omission in MT, we further propose selecting hard instances and utilizing GPT-4 to directly evaluate the performance of the models in mitigating these issues. We verify the rationality of these designed evaluation methods by experiments, followed by extensive results demonstrating the effectiveness of word alignment-based preference optimization to mitigate hallucination and omission. On the other hand, although it shows promise in mitigating hallucination and omission, the overall performance of MT in different language directions remains mixed, with slight increases in BLEU and decreases in COMET.
@@ 1 Introduction
Large language models (LLMs) have been evolving rapidly and showing predominant performance in many natural language processing (NLP) tasks (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). However, in machine translation (MT), the use of decoder-only LLMs is still limited due to issues such as model size (Xu et al., 2024a) and low-resource languages (Hendy et al., 2023). Conventional encoder-decoder MT models trained on parallel corpora still dominate in practice (Costa-jussà et al., 2022). One of the primary concerns of applying an LLM to MT is reliability. Although it does not happen frequently, an LLM is known to hallucinate (Dhuliawala et al., 2023; Zhang et al., 2023a; Bang et al., 2023) as it is pre-trained to predict the next token in very large-scale raw texts. Specifically in MT, LLM-based translation systems therefore could have the phenomena of hallucination and omission, which is also a long-term challenge in the field of MT (Yang et al., 2019; Vamvas and Sennrich, 2022), known as over- and under-translation. In particular, in the very recent WMT-2024 General Machine Translation Task Kocmi et al. (2024), a newly released LLM-based MT model Unbabel Tower Alves et al. (2024) has achieved the highest accuracy in most language pairs, demonstrating the promise of LLM in MT, but also showing the significance of the problem of hallucination and omission. As a result, we attempt to mitigate the hallucination and omission in LLM-based MT to improve its practicality in this work.
(a) Coverage distribution of different hallucination degree.
(b) Coverage distribution of different omission degree.
Figure 1: A preliminary experiment shows that higher coverage scores correlates to less hallucination and omission. The coverage scores are predicted by a word aligner Wu et al. (2023a). The human annotation of hallucination and omission is from HalOmi benchmark Dale et al. (2023b). Details about the dataset and word alignment model can be found in §5.1.
@@
Consequently, we propose Word Alignment Preference (WAP) that utilizes word alignment as a signal to optimize LLM-based MT models. WAP consists of three steps: diverse translation collection, preference data construction, and preference optimization. Specifically, we collect diverse translations with multiple existing translation tools, select chosen and rejected examples with the word aligner (Wu et al., 2023a), and optimize the model on preference data using direct preference optimization (DPO) (Rafailov et al., 2024).
Furthermore, the evaluation of hallucination and omission is challenging, and there is no existing evaluator specifically designed for this. Improving the BLEU and COMET score does not necessarily mean reducing hallucination and omission because there are other factors such as mistranslation and fluency. In addition, hallucination is relatively infrequent, although very severe once it does occur. Hence, to effectively evaluate it, we design extensive experiments that include testing on instances that potentially have the problem of hallucination and omission, and using GPT-4 as the evaluator with comprehensive analysis. Experimental analysis demonstrates the effectiveness of WAP in mitigating hallucination and omission in MT.
@@
  • As there is no particular benchmark for evaluating the performance of MT models on hallucination and omission. We design various experiments, including selecting hard instances and using LLM as an evaluator in §5.2. The effectiveness of the evaluation, as well as the proposed WAP has been validated through experiments and analysis in §6
@@ 2 Related work
Hallucination and omission in MT.
@@
Hallucinations are cases in which the model generates output that is partially or completely unrelated to the source sentence, while omissions are translations that do not include some of the input information (Dale et al., 2023b). Dale et al. (2023a) explore methods that leverage the internal workings of models and external tools, such as cross-lingual sentence similarity and natural language inference models, to detect and mitigate hallucinations in MT. HalOmi (Dale et al., 2023b) introduces an annotated dataset specifically designed to detect hallucinations and omissions. In Figure 1 and §5.2 we use HalOmi as a reference to assess how these two phenomena correlate to the coverage output of the GPT-4 evaluator and the word aligner, respectively. In particular, Yang et al. (2019) introduce the use of word alignment to reduce omission in MT, which partially inspires our idea.
Preference tuning for LLMs.
@@
LLMs are capable of completing tasks in the zero-shot or few-shot manner (Radford et al., 2019; Brown et al., 2020). In addition, performance in downstream tasks can also be enhanced by fine-tuning them with instruction datasets (Wei et al., 2022; Chung et al., 2024; Ouyang et al., 2022). However, acquiring instruction datasets is costly, while obtaining preferences for LLM responses is relatively easier (Rafailov et al., 2024). DPO (Rafailov et al., 2024) directly optimize LLM with preference data by removing an extra reward model. We utilize DPO in this work due to the ease of use and effectiveness. A contemporaneous preference-based method ALMA-R (Xu et al., 2024b), introduces contrastive preference optimization to fine-tune LLMs specifically using reference-free MT metrics and human annotation as preference. ALMA-R focuses on improving general LLM-based MT but we attempt to mitigate the hallucination and omission in MT. In addition, our preference data are made entirely automatically, which also draws the difference between ALMA-R and our work. The recently released LLM-based Unbabel Tower Alves et al. (2024) has achieved the best performance in most language pairs in WMT-2024 Kocmi et al. (2024), which may complement our findings in future work.
Word alignment.
Word-level information has been useful in many NLP tasks such as language pre-training Chi et al. (2021); Wu et al. (2021), cross-lingual sentence embedding Zhang et al. (2023b); Li et al. (2023); Miao et al. (2024), fine-grained visual language grounding Peng et al. (2023); Wu et al. (2023b, 2024), and particularly in word alignment for MT Bahdanau et al. (2015); Tu et al. (2016), which aligns the corresponding words in translations. Word aligners based on pre-trained language models Jalili Sabet et al. (2020); Dou and Neubig (2021); Nagata et al. (2020); Chousa et al. (2020) have outperformed previous ones based on statistical MT Och and Ney (2003); Dyer et al. (2013). WSPAlign (Wu et al., 2023a) is a pre-trained word aligner that outperforms most of the previous ones; hence we use it in the experiments.
@@
Figure 2: An illustration of WAP framework. The source is first translated by multiple MT tools, including human translation. An external word aligner is then utilized to predict the coverage score for each translation. Finally, translation with the highest and lowest coverage score are selected as preference pairs for preference optimization.
@@ 3.1 Gathering translation candidates
We start with the parallel training data in ALMA (Xu et al., 2024a). This parallel data encompasses five language pairs with human translations in both directions: $cs\leftrightarrow en$, $de\leftrightarrow en$, $is\leftrightarrow en$, $zh\leftrightarrow en$ and $ru\leftrightarrow en$. We employ ISO 639 language codes [^1] https://en.wikipedia.org/wiki/List_of_ISO_639_language_codes to denote languages. Specifically, “$cs$” corresponds to Czech, “$de$” to German, “$is$” to Icelandic, “$zh$” to Chinese and “$ru$” and “$en$” to Russian and English, respectively. To generate the translations we require, this dataset is translated in both directions using two well-known MT tools, including DeepL [^2] https://www.deepl.com/en/translator and ChatGPT (gpt-3.5-turbo-0613) [^3] https://platform.openai.com/docs/models/gpt-3-5-turbo. The prompt we use to translate sentences is shown in Figure 3. The original human-written translation in the training set is also utilized. In particular, Icelandic ($is$) is not supported by DeepL, therefore, we use Google Translate [^4] https://cloud.google.com/translate/docs/basic/translate-text-basic as an alternative.
Figure 3: The prompt for translating sentences.
@@ 3.2 Selecting chosen and rejected translation
After obtaining the translation candidates $(y^{1},...,y^{K})$, we use a state-of-the-art public word aligner, namely WSPAlign [^5] https://github.com/qiyuw/WSPAlign, to automatically annotate the degree of coverage for each translation. We follow the usage setting in the original paper of WSPAlign (Wu et al., 2023a). In particular, WSPAlign performs a bidirectional alignment and uses a threshold to filter out low-confident alignment of word pairs. Then, the ratio of the source words, that are aligned with at least one word, in the translation is taken as the coverage score, which will be used for the following preference annotation. The whole process to predict the coverage score is notated as $\mathrm{C}(\cdot,\cdot)$. Formally, the coverage score for a translation $y^{k}$ can be calculated by $\mathrm{C}(x,y^{k})\in[0.0,100.0]$. Subsequently, the preferred translation and the rejected translation are selected as follows:
@@ 3.3 Filtering
Note that the whole process of constructing the preference data is automatic, and the existing MT and word alignment models are not perfect. Even for human-annotated translation, quality is also an issue that cannot be ignored (Xu et al., 2024b), and can affect the performance of the model trained on it. Hence, noises are inevitable in both the translated texts and the preference choices. On the other hand, the MT tools we choose generally have good performance, it could happen that the generated translations are not diverse enough, leading to the preference signal being disrupted. To improve the quality of the constructed preference datasets as much as possible, multiple strategies are applied to filter out potential bad training instances:
@@
  • Remove the instance where the chosen and rejected translations are too semantically similar. Sentence embedding is a widely used technique for sentence similarity with low computation cost Gao et al. (2021); Wu et al. (2022); Xie et al. (2022); Zhao et al. (2024). In particular, LaBSE Feng et al. (2022) [^6] https://huggingface.co/sentence-transformers/LaBSE is used in our experiments. We notate it as $\mathrm{LB}(\cdot)$. The similarity threshold is empirically set as 0.9, i.e. $(x,y^{w},y^{l})$ is excluded from the dataset if $\mathrm{sim}(\mathrm{LB}(y^{w}),\mathrm{LB}(y^{w}))>0.9$. $\mathrm{sim}(\cdot,\cdot)\in[0.0,1.0]$ is cosine similarity.
@@
4 Details of dataset
Figure 4 presents the varying proportions of the “chosen” and “rejected” preference pairs from three sources: ChatGPT, DeepL, and Human. The figure indicates that most of the “chosen” translations originate from ChatGPT, while a significant portion of human-written translations are “rejected”. This observation supports the conclusion that human-written translations can also exhibit quality issues, as discussed in ALMA-R (Xu et al., 2024b). Examples in our constructed preference dataset are presented in §B.1.
@@ 4.1 Optimization LLM-based MT model
The final step is to optimize the LLM-based MT model on our preference data. Direct preference optimization (DPO) (Rafailov et al., 2024) is a simple but effective approach that directly optimizes the preference model on a pre-constructed static dataset. DPO has recently been applied to optimize LLM in preference data (Tunstall et al., 2023; Xu et al., 2024b) recently. We also utilize DPO as an optimization approach. Formally, the training objective is as follows,
@@
where $\sigma$ is the sigmoid function, $\pi$ is the model to be optimized, and $\pi_{ref}$ is the reference model. We use ALMA-13B [^8] https://github.com/fe1ixxu/ALMA as our base model, i.e., the starting point of $\pi$, in the experiments. ALMA-13B is also used as a reference model $\pi_{ref}$, but note that $\pi_{ref}$ will not be updated during training.
@@
Figure 5: Comparison of WAP and baseline in hard and easy instances. $N$ instances with the lowest COMET score by the baseline are selected from the test set as hard instances, and the remaining are easy instances. Results when $N=100$, $200$ and $500$ are presented. Refer to §C for the full numeric results of the entire test.
@@ 5.1 Baselines and evaluation datasets
We choose ALMA-13B [^9] https://huggingface.co/haoranxu/ALMA-13B as the baseline for all experiments in this paper, as well as the starting point of optimization. ALMA (Xu et al., 2024a) was trained from Llama Touvron et al. (2023) in two steps: initial fine-tuning on monolingual data and subsequent fine-tuning on a small set of high-quality parallel data. For fairly studying the effect of word alignment preference, we use the data used in the supervised fine-tuning in ALMA as the source dataset to construct our preference data in §3. Specifically, the source data was collected from WMT’17 (Bojar et al., 2017) to WMT’20 (Barrault et al., 2020), in addition to the development and text dataset from Flores-200 (Costa-jussà et al., 2022). After filtering, we finally make 20,074 and 2,226 preference triplets for training and development, respectively. For evaluation, the test set is from WMT22, except that $is\leftrightarrow en$ is from WMT21. The remaining data from WMT21 (except $is\leftrightarrow en$) is used as the development set. Specifically, 3485, 4021, 2000, 3912, 4053 examples are included in the test set for $cs\leftrightarrow en$, $de\leftrightarrow en$, $is\leftrightarrow en$, $zh\leftrightarrow en$, and $ru\leftrightarrow en$, respectively. The detailed experimental setup is introduced in §A.
@@
In particular, we want to validate whether our proposed method is capable of mitigating hallucination and omission in MT. Hence, we also use HalOmi (Dale et al., 2023b) in the experiments. HalOmi is an evaluation benchmark for the detection of hallucination and omission in MT. It contains fine-grained sentence-level and token-level annotations of full and partial hallucinations and omissions that cover 18 language directions. Each instance in the data set was annotated in “No hallucination and omission”, “Small hallucination and omission”, “Partial hallucination and omission” or “Full hallucination and omission” by humans. In this paper, we use it to test the performance of GPT-4 as an evaluator. Details are in §5.2.